OpenStudy (anonymous):

solve 15t²-15t-30=0 medal awarded

OpenStudy (anonymous):

dont i find factors for -thirty

OpenStudy (amistre64):

factor out the 15 first ... then you have the usual: 15 (x^2+ ...)

OpenStudy (amistre64):

you want the first term to be "nekkid"

OpenStudy (anonymous):

nekkid do you mean naked??

OpenStudy (dan815):

like a catt

OpenStudy (dan815):

no like an apple

OpenStudy (amistre64):

:)

OpenStudy (amistre64):

that 15 is just dressing, we need to take it off .....

OpenStudy (dan815):

how dare you not simplify this equation

OpenStudy (anonymous):

get off of here please i dont like your sarcasm any way amistre ok so i forgot how to strip the t away from the 15

OpenStudy (amistre64):

essentially, you divide everything by 15

OpenStudy (amistre64):

15t^2/15 - 15t/15 - 30/15 = 0/15 t^2 - t - 2 = 0

OpenStudy (anonymous):

okay so that gives me t²-t-2=0

OpenStudy (amistre64):

correct, and that should be more than simple enough to determine solutions for :)

OpenStudy (anonymous):

well i can add two to both sides and solve that would leave t²-t=2 so ill solve for t and t² t² would be 2 and t would be 2

OpenStudy (amistre64):

m+n = -1 mn = -2 factors of -2 are: 1 -2 -2+1 = -1 ; so lets reverse signs: 1,-2 becomes -1,2

OpenStudy (anonymous):

-1 and 2

OpenStudy (amistre64):

id stick with those :)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

okay x²-8x+15=0

OpenStudy (amistre64):

m+n = -8 mn = 15 factors of 15 are?

OpenStudy (anonymous):

1 and 15 5 and 3

OpenStudy (amistre64):

then lets use 5 and 3 in this fashion: -5-3 = -8 (-5)(-3) = 15 those work marvelously so reverse the signs: -3,-5 becomes 3,5

OpenStudy (anonymous):

i think im starting to get this

OpenStudy (anonymous):

okay somethi like this (x-2)²=49

OpenStudy (amistre64):

well, hopefully we know that 7^2 = 49; and (-7)^2 = 49 x-2 = 7 or x-2 = -7

OpenStudy (amistre64):

if that wasnt a perfect square of there, the concept is the same: (x-r)^2 = k when x-r = +- sqrt(k); x = r +- sqrt(k)

OpenStudy (anonymous):

yes i know that

OpenStudy (anonymous):

okay i still need help...